A light on the ground at Lilys building is 30meters away from the building. Jade is 2 meters tall. She walks from the light to the building at 1 meters per second. How fast is the shadow of Jade on the building changing when she is 15 meters from the building.


1
Expert's answer
2021-07-01T16:12:10-0400



"\text{By using similar triangle property,}\newline
\frac{2}{x}=\frac{h}{30}\newline
h=\frac{60}{x}\newline
\text{Differentiate wrt x,}\\
\frac{dh}{dx}=\frac{-60}{x^2}\\
Given, \frac{dx}{dt}=1m/sec.\newline
\text{
we need to find} \frac{dh}{dt}.\newline
Therefore, \frac{dh}{dt}=\frac{dh}{dx}×
\frac{dx}{dt}\newline
=\frac{-60}{x^2}×1\newline
=\frac{-60}{x^2}\newline
\text{
Given, when she is 15 m from the building.} \newline
\implies x=30-15=15m\newline
Therefore,\newline
\frac{dh}{dt}=\frac{-60}{15^2}=\frac{-4}{15}=-0.267m/sec"